Selecting Players How many ways can 4 baseball


players and 3 basketball players be selected from


12 baseball players and 9 basketball players


1
Expert's answer
2022-06-13T16:12:56-0400
"\dbinom{12}{4}\dbinom{9}{3}=\dfrac{12!}{4!(12-4)!}\cdot \dfrac{9!}{3!(9-3)!}"

"=\dfrac{12(11)(10)(9)}{1(2)(3)(4)}\cdot\dfrac{9(8)(7)}{1(2)(3)}"

"=495\cdot84=41580"